OpenStudy (anonymous):

y=x^2-7x+12 need axis of symmetry, vertex, mx/min value, y-intercept HELP PLEASEEE AND THANKS!!!

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case its \[-\frac{-7}{2\times 1}=\frac{7}{2}\]

OpenStudy (anonymous):

that is also your axis of symmetry, \[x=\frac{7}{2}\]

OpenStudy (anonymous):

okay i just took note of that equation to remember..could you help me out with vertex also? @satellite73